A circle has radius 3 and it's centre lie on the line y=x-1. Fi d the equation of circle, if it passes through (7,3)...

Grade:11

1 Answers

Arun
5 years ago
 
Dear Ishita
 
let centres are (h,k)
 
(x-h)^2 + (y-k)^2 = 3^2
 
also k = h-1
 
then pqss it through (7,3)
 
you will get the answer.